OpenStudy (s):

how to find x in x^2-x-1 ?

OpenStudy (radar):

As is, you do not have an equation. What is that expression equal to?

OpenStudy (s):

to 0

OpenStudy (lgbasallote):

 i found x

OpenStudy (radar):

The quadratic formula seems like a good candiate. OMG lgbasallote has already found x!

OpenStudy (lgbasallote):

teachers should phrase it as solve for x :P find x is tooo easy (unless x happens to be wally)

OpenStudy (s):

that was not helpful at all , but I understand your point

OpenStudy (lgbasallote):

on a helpful note though...use either completing the square or q.f.

OpenStudy (lgbasallote):

it is not factorable...i can give you that

OpenStudy (radar):

\[x ^{2}+x-1=0\]A=1 B=1 C=-1 Use quadratic formula.

OpenStudy (anonymous):

C would be +1, when you move it to the other side of the equation

OpenStudy (radar):

x=[-1+/-sqrt(1+4]/2 \[=(-1\pm \sqrt{5})/2\]